OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations.

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

Elimination :)

OpenStudy (texaschic101):

2x + y = 7 5x + y = 9 --->(-1) ------------- 2x + y = 7 -5x - y = -9 (result of multiplying by -1) ------------add -3x = -2 x = 2/3 now sub 2/3 in for x in either equation to find y 2x + y = 7 2(2/3) + y = 7 4/3 + y = 7 y = 7 - 4/3 y = 21/3 - 4/3 y = 17/3 check.. 5x + y = 9 5(2/3) + 17/3 = 9 10/3 + 17/3 = 9 27/3 = 9 9 = 9 (correct) x = 2/3 and y = 17/3 I always hated the ones with fractions...lol

OpenStudy (anonymous):

What if it's a problem like this? Find the solution to the equations. 2x - y = -3 x + y = 0